# Generate, Transduct, Adapt: Iterative Transduction with VLMs

### v1

@@ Abstract
Transductive zero-shot learning with vision-language models leverages image-image similarities within the dataset to achieve better classification accuracy compared to the inductive setting. However, there is little work that explores the structure of the language space in this context. We propose GTA-CLIP, a novel technique that incorporates supervision from language models for joint transduction in language and vision spaces. Our approach is iterative and consists of three steps: (i) incrementally exploring the attribute space by querying language models, (ii) an attribute-augmented transductive inference procedure, and (iii) fine-tuning the language and vision encoders based on inferred labels within the dataset. Through experiments with CLIP encoders, we demonstrate that GTA-CLIP, yields an average performance improvement of 8.6% and 3.7% across 12 datasets and 3 encoders, over CLIP and transductive CLIP respectively in the zero-shot setting. We also observe similar improvements in a few-shot setting. We present ablation studies that demonstrate the value of each step and visualize how the vision and language spaces evolve over iterations driven by the transductive learning.
Figure 1: Overview of GTA-CLIP. (a) Vision-language models (VLMs) such as CLIP [36] enable zero-shot classification using similarity between text embeddings of class prompts and images. (b) Transductive CLIP [61] exploits the structure of entire image dataset to assign images to classes improving accuracy. (c) Our approach, GTA-CLIP, iteratively (i) induces structure over the classes in language space by generating attributes driven by the pairwise confusions, (ii) performing attribute-augmented transductive inference, and (iii) adapting CLIP encoders using the inferred labels. (d) Across 12 datasets we improve upon CLIP and transductive CLIP by 8.6% and 4.0% using VIT-B/32, and similarly for other encoders. Significant improvements are also reported in the few-shot setting.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have enabled zero-shot image understanding across a wide range of domains. A common approach assigns images to classes based on the similarity between image and text embeddings generated by models like CLIP [36], and is a basis for many zero-shot methods in image classification [66, 67, 63], segmentation [54, 37, 20, 21], and detection [27, 65] (Fig. 1a). However, in practical scenarios, the images requiring labeling are often available in advance. For example, an ecologist may have access to a large collection of animal images they aim to categorize by species. In such cases, transductive inference is better suited, as it leverages the dataset’s inherent structure to refine model predictions (Fig. 1b).
However, few approaches have fully explored the structure of the label space derived from language in these scenarios. For instance, linking similar descriptions or attributes can produce more coherent text-based prototypes for each class. Furthermore, attributes and improved alignment with images can be used to adapt vision and language models to the specific dataset. This has the potential to improve tranductive learning with CLIP, when deployed on novel domains or fine-grained recognition tasks.
@@
This motivates our approach, GTA-CLIP, we explore transductive learning by leveraging structure in both the language and vision spaces (Fig. 1c and Alg. 1). Our method begins by querying a language model to populate the language space: we start with an initial set of attributes for each category and then dynamically expand this space by generating discriminative attributes driven by pairwise confusion between classes. This strategy maintains tractability while improving performance. We design a transductive inference procedure for vision-language models (VLMs) augmented by these attributes. Finally, we use the inferred labels and attributes to adapt the underlying VLM to the target dataset. This cycle of generation, transduction, and adaptation–hence the name GTA–repeats over several iterations, progressively improving recognition performance.
@@
We present experiments on a benchmark of 12 datasets using various CLIP encoders, where our approach achieves 8.6% improvement over CLIP and 3.7% improvement over the current state-of-the-art transductive CLIP [61] on average (Fig. 1d and Table 1). Ablation studies show that each component of our method contributes to these performance gains. For instance, while attribute-augmented transduction improves performance on average, it is not consistently effective without model fine-tuning. Similarly, incrementally expanding the attribute space provides improvements in fine-grained domains while keeping learning tractable.
We also report results for few-shot setting (Table 2) and a zero-shot setting where labels from different categories within the same domain are available during training (Table 3). In both cases, GTA-CLIP significantly outperforms transductive CLIP [61] and prior approaches. Additionally, we analyze how the language and vision spaces evolve over iterations providing insights into the performance improvements. Our experiments indicate that gains from attribute generation, transduction, and model adaptation are complementary to traditional labeling efforts, which is of practical value as it provides practitioners different avenues to improve labeling accuracy on their datasets. Code will be released at https://github.com/cvl-umass/GTA-CLIP.
@@ 2 Related Work
Transductive learning [46] is suited to scenarios where a model’s predictions need to be accurate on a specific dataset instead of future unseen data. Approaches assume access to both labeled training data and unlabeled test data during training. Techniques include label propagation [10], learning class prototypes by clustering [48, 59], assignment using optimal transport [50], and others [4, 19]. The setting is similar to semi-supervised learning, where techniques like pseudo-labeling [62, 6, 3], entropy minimization [13], and self-training [56, 53] have proven effective.
@@
Zero-shot transduction has been investigated previously using image generation [12, 49] and attribute-based approaches [55, 60]. Recent approaches use CLIP to facilitate zero-shot learning without domain-specific training. For example, ZLaP [16] improves CLIP through label propagation, while [25] iteratively estimate assignments and class prototypes guided by language. TransCLIP (NeurIPS’24) [61] presents an efficient approach for large-scale transduction, employing a block majorization-minimization (BMM) algorithm [15, 38] to optimize an objective comprising: a Gaussian mixture model, a Laplacian regularizer, and a KL divergence term that aligns assignments with image-text probabilities across the dataset. We adapt this formulation to our setting as it represents the current state-of-the-art and is scalable.
@@
Improving Zero-shot with Attributes. Previous work has used large language models (LLMs) to expand the attribute space and improve zero-shot classification. For instance, [35, 26, 29] use LLMs to augment category descriptions beyond simple class names (e.g., describing a tiger as having stripes and claws) to improve zero-shot classification and interpretability in CLIP-based models. Rather than relying solely on language models, other approaches aim to identify a concise set of recognition attributes [57, 7]. In this work, we adopt a similar approach by querying language models like GPT [1] and Llama [45] to populate the attribute space. Furthermore, we dynamically expand this space by adding attributes to classes that are frequently confused, enabling better discrimination.
@@
Adapting CLIP. Prior work has shown that augmenting CLIP with attributes does not improve zero-shot recognition, especially when deployed on out-of-domain or fine-grained datasets. In this case, model adaption is necessary. Techniques range from learning language and vision prompts [67, 66], adding learnable layers [11, 63], to full fine-tuning [42, 44, 51, 64]. A different line of work addresses fine-tuning without paired image and text data. WiSE-FT [51] and LaFTer [28] use ensembles, while others [18, 24, 40] show the value of large-scale fine-tuning with image-text data aligned at the category level. We adopt the approach of AdaptZSCLIP [40] which stochastically pairs images with attributes within a category and modifies the CLIP objective to account for the weaker supervision. However, unlike prior work that rely on labeled examples, we perform adaption in a zero-shot manner.
In summary, our approach uses language models to iteratively discover attributes, and improve transduction and enable fine-tuning in a zero-shot manner. While these ideas have been explored individually, the combination is novel, and we demonstrate to significant gains over CLIP, transductive CLIP, zero-shot classification with attributes, and fine-tuning on target domains, across a range of benchmarks. For example, no prior work has shown fine-tuning is effective without any labeled data in the target domain. The benefits also observed in the few-shot setting.
@@ 3 Methodology
The input to our approach is a set of images ${\cal X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ and a set of classes ${\cal Y}=\{y_{i}\}_{i=1}^{M}$. In the zero-shot setting the goal is to assign each image to one of the $M$ classes. In the few-shot setting we are also provided with a few labeled examples ${\cal D}_{train}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{K}$ with $x\in{\cal X}_{train}$, ${\cal X}_{train}\cap{\cal X}=\O$, and $y\in{\cal Y}$. We also consider a setting where labeled data comes from a different set of classes, i.e., $y\in{\cal Y}_{train}$ where ${\cal Y}_{train}\cap{\cal Y}=\O$. This setup is used in approaches where labeled data from a set of base categories is used to adapt CLIP on the target domain. We report the mean per-class accuracy on the target set of images ${\cal X}$ given their ground-truth labels. To enable zero-shot learning we assume an image encoder $\theta(\cdot)$ and a text encoder $\phi(\cdot)$ such that $\theta(\mathbf{x})^{\top}\phi(\mathbf{y})$ is high for image $\mathbf{x}$ and text $\mathbf{y}$ pairs that are similar. We experiment with a variety of encoder pairs based on CLIP. In addition we assume access to a language model (e.g., Llama3) which we can query to generate attributes for each class.
@@ 3.1 GTA-CLIP formulation
GTA-CLIP maintains a list of attributes indexed by class denoted by ${\cal A}=({\cal A}_{j})_{j=1}^{M}$, where ${\cal A}_{j}=\{\mathbf{a}_{j,k}\}_{k=1}^{n_{i}}$ denotes the set of text attributes for the class $j$. The number of attributes $n_{j}$ can vary across classes. Following the TransCLIP [61] formulation we maintain $\boldsymbol{\mu}=(\boldsymbol{\mu}_{j})_{j=1}^{M}$ and $\boldsymbol{\Sigma}=(\boldsymbol{\Sigma}_{j})_{j=1}^{M}$ denoting the Gaussian mixture model (GMM) mean and diagnonal variance for each class. In addition we maintain a matrix of softmax class assignments $\mathbf{z}\in[0,1]^{N\times M}$, where $N$ is the number of query images and $M$ is the number of classes. In other words $\mathbf{z}_{i,\cdot}\in\Delta_{M}$ reflects the probability of assignment over all the classes, where $\Delta_{M}$ is the $M$-dimensional probability simplex. Given a class $j\in\mathcal{Y}$ the vertical slices $\mathbf{z}_{\cdot,j}\in[0,1]^{N}$ represents the probability that a specific query image belongs to class $j$. After inference the class label for each image $i$ can be obtained as $\operatorname{argmax}_{j}\mathbf{z}_{i,j}$.
@@
6:     $\mathcal{A}$leftarrow GenerateAttributes($\mathcal{Y},\mathcal{A},\theta,\phi$)
@@
8:     $\mathbf{z},\boldsymbol{\mu},\boldsymbol{\Sigma}$leftarrow Transduct($\mathcal{X},\mathcal{Y},\mathcal{A},\theta,\phi$)
@@
10:     $\theta,\phi$leftarrow Adapt($\mathcal{X},\mathcal{Y},\mathbf{z},\theta,\phi$)
@@
The second term is a Laplacian regularizer commonly seen in spectral clustering [30, 41] and semi-supervised learning settings [2, 58]. Here $w_{i,j}$ denotes the affinity between images $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, and this term encourages images with high affinity to have similar predictions $\mathbf{z}$. Following TransCLIP, we set $w_{i,j}=\max(0,\mathbf{f}_{i}^{\top}\mathbf{f}_{j})$ resulting in a positive semi-definite affinity matrix $\mathbf{W}=[w_{i,j}]$ and faster optimization procedure due to a convex relaxation.
@@
[…]
@@ 3.2 Optimization
where [domain] is coarse category, e.g. “birds” for CUB200 [47], [class] is the common name of the category, and [attribute] is a specific attribute. For example, one such description is “A bird with a small, round body shape, indicative of a Baird’s Sparrow.”
@@
To keep this tractable we only generate attributes for the most confused classes. We first update $\mathbf{z}$ by running attribute-augmented transductive inference given the current model and set of attributes $\mathcal{A}$ (Step 2) [^1] Since transductive inference is relatively efficient, we find it beneficial to run this step before invoking the more expensive generate step (see § 8). Then, we find the images $\mathbf{x}_{i}$ for which the difference in the top $2$ probabilities in $\mathbf{z}_{i,\cdot}$ is lower than a threshold of $\alpha$:
[…]
@@ 2. Attribute-Augmented Transductive Inference
We finally fine-tune CLIP encoders $\theta,\phi$ using the current set of attributes $\mathcal{A}$ and the inferred labels $\mathbf{z}$. For each class $j$ we find the top $k$ images with the highest scores based on $\mathbf{z}_{\cdot,j}$. The set of images and corresponding attributes provide a coarse form of supervision for fine-tuning. Specifically, we use the modified objective of AdaptCLIPZS [40] which takes into account class-level supervision and false negative associates since multiple text-image pairs can be considered correctly aligned in a single mini-batch training. For the few-shot setting, we simply include the labeled examples to our samples.

### v2

@@ Abstract
Transductive zero-shot learning with vision-language models leverages image-image similarities within the dataset to achieve better classification accuracy compared to the inductive setting. However, there is little work that explores the structure of the language space in this context. We propose GTA-CLIP, a novel technique that incorporates supervision from language models for joint transduction in language and vision spaces. Our approach is iterative and consists of three steps: (i) incrementally exploring the attribute space by querying language models, (ii) an attribute-augmented transductive inference procedure, and (iii) fine-tuning the language and vision encoders based on inferred labels within the dataset. Through experiments with CLIP encoders, we demonstrate that GTA-CLIP yields an average performance improvement of 9.5% and 4.0% across 12 datasets and 3 encoders, over CLIP and transductive CLIP respectively in the zero-shot setting. We also observe similar improvements in a few-shot setting. We present ablation studies that demonstrate the value of each step and visualize how the vision and language spaces evolve over iterations driven by the transductive learning.
Figure 1: Overview of GTA-CLIP. (a) Vision-language models (VLMs) such as CLIP [36] enable zero-shot classification using similarity between text embeddings of class prompts and images. (b) Transductive CLIP [61] exploits the structure of the entire image dataset to assign images to classes improving accuracy. (c) Our approach, GTA-CLIP, iteratively (i) induces structure over the classes in language space by generating attributes driven by the pairwise confusions, (ii) performing attribute-augmented transductive inference, and (iii) adapting CLIP encoders using the inferred labels. (d) Across 12 datasets we improve upon CLIP and transductive CLIP by 9.5% and 4.0% using VIT-B/16, and similarly for other encoders. Significant improvements are also reported in the few-shot setting.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have enabled zero-shot image classification across diverse domains. These models, such as CLIP [36], assign images to classes based on the similarity between image and text embeddings, forming the basis of various zero-shot approaches in classification [66, 67, 63], segmentation [54, 37, 20, 21], and detection [27, 65] (Fig. 1a). However, in many practical scenarios, the images requiring classification are known in advance. For example, an ecologist might have a large collection of animal images that need to be categorized by species. In such cases, transductive inference is more suitable, as it leverages the dataset’s inherent structure to refine predictions (Fig.1b).
Despite the success of transductive inference with VLMs, existing approaches often overlook the rich structure of the label space derived from language. For instance, linking semantically similar descriptions or attributes can yield more coherent class prototypes, while aligning these attributes with image features can enable model adaptation on the specific dataset. This strategy can be advantageous for zero-shot and few-shot recognition, especially in novel or fine-grained domains where labeled data is scarce.
@@
To address this gap, we propose GTA-CLIP, a transductive learning approach that exploits structure in both the language and vision spaces (Fig.1c and Alg.1). Our method begins by querying a language model to populate the language space: starting with an initial set of attributes per category, we dynamically expand this space by generating discriminative attributes based on pairwise confusion between classes. This strategy improves class separation while maintaining computational tractability. We then design a transductive inference procedure that refines predictions using these attributes. Finally, we adapt the underlying VLM to the target dataset using inferred labels and attributes. This iterative cycle of Generation, Transduction, and Adaptation—hence the name GTA—progressively improves recognition performance.
@@
We present experiments on a benchmark of 12 datasets using various CLIP encoders, where our approach achieves 8.6% improvement over CLIP and 3.7% improvement over the current state-of-the-art transductive CLIP [61] on average (Fig. 1d and Table 1). Notably, on a dataset like CUB with about 12k images, the whole process completes in 12-20 minutes on a single A100 GPU (see § 5.6). Ablation studies demonstrate that each component of our method contributes to these gains. Specifically, while attribute-augmented transduction improves performance on average, it is most effective when paired with model fine-tuning. Similarly, dynamically expanding the attribute space benefits fine-grained domains while keeping learning efficient. We visualize how the language and vision spaces evolve over iterations providing insights into the performance improvements. Further, we demonstrate GTA-CLIP’s advantages in a few-shot setting (Table 2) and a zero-shot setting where labeled examples from related categories are available during training (Table 3). In both cases, our approach outperforms transductive CLIP [61] and prior methods.
To summarize, our main contribution is to demonstrate that zero- and few-shot classification can be significantly improved in a transductive setting by integrating attribute generation, transductive inference, and model adaptation into a unified framework. While prior work has explored these components in isolation, to the best of our knowledge, this is the first work to show that their benefits are complementary and can be effectively leveraged in label-scarce scenarios. Our approach is of practical value as it provides end users with another avenue to improve labeling accuracy on their target dataset, alongside traditional labeling efforts. Our code is released at https://github.com/cvl-umass/GTA-CLIP.
@@ 2 Related Work
Transductive learning [46] is well-suited for scenarios where a model’s predictions must be accurate on a specific dataset rather than on unseen future data. Access to the entire unlabeled test set enables inference through methods such as label propagation [10, 50], clustering [48, 59, 50], among others [4, 19]. This setting closely resembles semi-supervised learning, where techniques like pseudo-labeling [62, 6, 3], entropy minimization [13], and self-training [56, 53] have proven effective.
@@
Zero-shot transduction has been previously explored using image generation [12, 49] and attribute-based approaches [55, 60], while more recent methods leverage VLMs to estimate initial class prototypes from language. For example, ZLaP [16] improves CLIP through label propagation, while [25] iteratively estimate assignments and class prototypes. TransCLIP [61] presents an efficient approach for large-scale zero-shot transduction, employing a block majorization-minimization (BMM) algorithm [15, 38] to optimize an objective comprising: a Gaussian mixture model, a Laplacian regularizer, and a KL divergence term that aligns assignments with image-text probabilities across the dataset. We extend this state-of-the-art by incorporating class-specific attributes in the KL divergence term, enabling better alignment of image features with the semantic structure of the dataset.
@@
Improving Zero-shot with Attributes. Large language models (LLMs) have been used to improve zero-shot classification by expanding attribute spaces beyond simple class names. For example, [35, 26, 29] employ LLMs to generate rich category descriptions (e.g., describing a tiger as having stripes and claws) to improve both classification accuracy and interpretability in CLIP-based models. Beyond LLM-based augmentation, other approaches focus on identifying a concise and discriminative set of attributes for recognition [57, 7]. Inspired by both strategies, we leverage language models such as GPT [1] and LLaMA [45] to initially populate the attribute space. However, rather than relying on static expansions, we introduce a dynamic refinement process: attributes are iteratively added to classes that are frequently confused, improving class separability.
@@
Adapting CLIP. Prior work has shown that augmenting CLIP with attributes does not significantly improve zero-shot recognition, particularly in out-of-domain or fine-grained datasets. In such cases, model adaptation is necessary. Existing techniques range from learning language and vision prompts [67, 66] to incorporating learnable layers [11, 63] or performing full fine-tuning [42, 44, 51, 64]. A different line of work addresses fine-tuning without paired image and text data. WiSE-FT [51] and LaFTer [28] use ensembles, while others [18, 24, 40] show the value of large-scale fine-tuning with image-text data aligned at the category level. We build on the approach of AdaptCLIPZS [40], which stochastically pairs images with attributes within a category and modifies the CLIP objective to accommodate weaker supervision. However, while all the above approaches rely on labeled examples, such as images with class labels on the target domain, our method enables adaptation without any annotated data.
Our key contribution is unifying attribute generation, model adaptation, and transductive inference within a single framework for zero- and few-shot classification. Iterative and stage-wise learning can be viewed as the optimization of a single objective, enabling both label inference within the dataset and end-to-end fine-tuning of the underlying VLMs on target domains (Alg. 1). While these ideas have been explored individually, their integration is novel and leads to significant improvements over the current state-of-the-art across diverse datasets, with minimal additional computational requirements.
@@ 3 Methodology
The input to our approach is a set of images ${\cal X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ and a set of classes ${\cal Y}=\{y_{i}\}_{i=1}^{M}$. In the zero-shot setting the goal is to assign each image to one of the $M$ classes. In the few-shot setting we are also provided with a few labeled examples ${\cal D}_{train}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{K}$ with $\mathbf{x}\in{\cal X}_{train}$, ${\cal X}_{train}\cap{\cal X}=\O$ , and $y\in{\cal Y}$.
We also consider a setting where labeled data comes from a different set of classes, i.e., $y\in{\cal Y}_{train}$ where ${\cal Y}_{train}\cap{\cal Y}=\O$ . This setup is used in approaches where labeled data from a set of base categories is used to adapt CLIP on the target domain.
We report the mean per-class accuracy on the target set of images $\cal X$} given their ground-truth labels. To enable zero-shot learning we assume an image encoder $\theta(\cdot)$ and a text encoder $\phi(\cdot)$ such that $\theta(\mathbf{x})^{\top}\phi(\mathbf{y})$ is high for image $\mathbf{x}$ and text $\mathbf{y}$ pairs that are similar. We experiment with a variety of encoder pairs based on CLIP framework. In addition we assume access to a language model (e.g., Llama3 or GPT-4o) which we can query to generate attributes for each class.
@@ 3.1 GTA-CLIP formulation
GTA-CLIP maintains a list of attributes indexed by class denoted by ${\cal A}=({\cal A}_{j})_{j=1}^{M}$, where ${\cal A}_{j}=\{\mathbf{a}_{j,k}\}_{k=1}^{n_{j}}$ denotes the set of text attributes for the class $j$. The number of attributes $n_{j}$ can vary across classes. Like the TransCLIP [61] formulation we maintain $\boldsymbol{\mu}=(\boldsymbol{\mu}_{j})_{j=1}^{M}$ and $\boldsymbol{\Sigma}=(\boldsymbol{\Sigma}_{j})_{j=1}^{M}$ denoting the Gaussian mixture model (GMM) mean and diagnonal variance for each class.
In addition we maintain a matrix of softmax class assignments $\mathbf{z}\in[0,1]^{N\times M}$, where $N$ is the number of query images and $M$ is the number of classes. In other words $\mathbf{z}_{i,\cdot}\in\Delta_{M}$ reflects the probability of assignment over all the classes, where $\Delta_{M}$ is the $M$-dimensional probability simplex. Given a class $j\in\mathcal{Y}$ the vertical slices $\mathbf{z}_{\cdot,j}\in[0,1]^{N}$ represents the probability that a specific query image belongs to class $j$. After inference the class label for each image $i$ can be obtained as $\operatorname{argmax}_{j}\mathbf{z}_{i,j}$.
@@
6:  $\mathcal{A}$ $\leftarrow$ GenerateAttributes($\mathcal{Y},\mathcal{A},\theta,\phi$)
@@
8:  $\mathbf{z},\boldsymbol{\mu},\boldsymbol{\Sigma}$ $\leftarrow$ Transduct($\mathcal{X},\mathcal{Y},\mathcal{A},\theta,\phi$)
@@
10:  $\theta,\phi$ $\leftarrow$ Adapt($\mathcal{X},\mathcal{Y},\mathbf{z},\theta,\phi$)
@@
The second term is a Laplacian regularizer commonly seen in spectral clustering [30, 41] and semi-supervised learning settings [2, 58]. Here $w_{i,j}$ denotes the affinity between images $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, and this term encourages images with high affinity to have similar predictions $\mathbf{z}$. We set $w_{i,j}=\max(0,\mathbf{f}_{i}^{\top}\mathbf{f}_{j})$ resulting in a positive semi-definite affinity matrix $\mathbf{W}=[w_{i,j}]$ and faster optimization procedure due to a convex relaxation.
@@
$\hat{\mathbf{y}}_{i,j}=\frac{\exp(\bar{s}_{i,j})}{\sum_{j=1}^{M}\exp(\bar{s}_{i,j})},\text{~where~}\bar{s}_{i,j}=\frac{1}{n_{j}}\sum_{k=1}^{n_{j}}\theta(\mathbf{x}_{i})\phi(\mathbf{a}_{j,k}).$ (4)
@@ 3.2 Optimization
where [domain] is coarse category, e.g. “birds” for CUB [47], [class] is the common name of the category, and [attribute] is a specific attribute. For example, one such description is “A bird with a small, round body shape, indicative of a Baird’s Sparrow.”
@@
To keep this tractable we only generate attributes for the most confused classes. We first update $\mathbf{z}$ by running attribute-augmented transductive inference given the current model and set of attributes $\mathcal{A}$ (Step 2) [^1] We find it beneficial to run the transductive step before invoking the generate step (see Appendix Table 9). Then, we find the images $\mathbf{x}_{i}$ for which the difference in the top $2$ probabilities in $\mathbf{z}_{i,\cdot}$ is lower than a threshold of $\alpha$:
$\mathcal{CC}=\{(i,\{c_{1},c_{2}\})\hskip 2.84544pt|\hskip 2.84544pt\mathbf{z}_{i,c_{1}}-\mathbf{z}_{i,c_{2}}\leq\alpha;c1<c2\}.$
@@ 2. Attribute-Augmented Transductive Inference
We finally fine-tune CLIP encoders $\theta,\phi$ using the current set of attributes $\mathcal{A}$ and the inferred labels $\mathbf{z}$. For each class $j$ we find the top $k$ images with the highest scores based on $\mathbf{z}_{\cdot,j}$. The set of images and corresponding attributes provide a coarse form of supervision for fine-tuning. Specifically, we adopt the objective of AdaptCLIPZS [40] which takes into account class-level supervision and false negative associates since multiple text-image pairs can be considered correctly aligned in a single mini-batch training. For the few-shot setting, we simply include the labeled examples to our samples.